What are the components of benzene? Specify the percentage of each.
1
Expert's answer
2017-03-14T02:42:06-0400
Benze has a formula C6H6.
Components are carbon (C) and hydrogen (H). Molecular weight is:
M=6×A(C)+6×A(H)
M=6×12.01+6×1.01=78.11
%(C)=(6×12.01)/78.11×100%=92.26 %
%(H)=(6×1.01)/78.11×100%=7.74 %